What is the average speed if a car covers 200 miles in 2 hours and 30 minutes

Answer by TimothyLamb(4379)

---
s = d/t
s = 200/(2+(30/60))
s = 200/(2+(1/2))
s = 200/((4/2)+(1/2))
s = 200/(5/2)
s = 400/5
s = 80 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php